CAN SOMEONE HELP ME WITH THIS PROBLEM WHEN I CHECK IT THE ANSWER DOESN'T MATCH:
4W+9-W+4=3W+11
4W-W+13=3W+11
3W+13=3W+11
3W=24
W=8

CHECK:
4(8)+9-8+4=3(8)+11
32+9-8+4=24+11
41-4=35
37=35
WHAT AM I DOING WRONG?

Answer by ilana(307)

Found your mistake: you can't go from 3W+13=3W+11 to 3W=24. Instead, from 3W+13=3W+11, you could subtract 3W from both sides and get 13=11, showing you there is no solution to this equation. You could also subtract 11 from both sides, giving you 3W+2=3W. Then subtract 3W from both sides and get 2=0, again showing there is no solution. Regardless, this equation has no solution.